at 25 degrees celsius the rate constant for the first order decomposition of a pesticide solution is 6.40x10^-3 min^-1. If the starting concentration of pesticide is 0.0314 M, what concentration will remain after 62.0 min @ 25 degrees celsius?

Correct answer is 2.11X10^-2M
Can someone please explain how to get this answer?
Answer #1

K   = 2.303/t log[A0]/[A]

K   = 6.4*10^-3 min^-1

[A0] = 0.0314M

t          = 64 min

K   = 2.303/t log[A0]/[A]

6.4*10^-3 = 2.303/64 log0.031/[A]

log0.031/[A]   = 6.4*10^-3 *64/2.303

log0.031/[A] = 0.1778

0.031/[A]      = 10^0.1778

0.031/[A]    = 1.5059

[A]               = 0.031/1.5059   = 0.0211M = 2.11*10^-2 M >>>>answer